OpenStudy (anonymous):

Verify the identity cos(2x) / sin(x) cos(x) = cot(x) - tan(x)

OpenStudy (loser66):

go backward from RHS cot x - tan x =\[\frac{ cosx }{ sinx }- \frac{ sinx }{ cosx }= \frac{ \cos^2 - \sin^2 }{ sinxcox }=\frac{ \cos2x }{ sinxcosx }\]

OpenStudy (anonymous):